I am asking for someone to help me solve this problem, 3x - 5(3x - 7) = 2(x + 9) + 45. Please help me.
Answer by checkley71(8403)

3x-5(3x-7)=2(x+9)+45 or 3x-15x+35=2x+18+45 or 3x-15x-2x=18+45-35 or
-14x=28 or x=-2 IS THE ONLY ANSWER.
proof 3*-2-5(3*-2-7)=2(-2+9)+45 or -6-5(-6-7)=2(7)+45 or -6-5(-13)=14+45 or
-6+65=59 or 59=59 THESE ARE THE STEPS FOR THE PROOF.